OpenStudy (anonymous):

How do you solve for x. x^3 + x +4 = 0

OpenStudy (anonymous):

x(x^2+1) = -4

OpenStudy (anonymous):

x(x^2 + 1) = -2*2

OpenStudy (anonymous):

so whats the answer?

OpenStudy (anonymous):

it will be either x = -2 or, x^2 +1 = 2

OpenStudy (anonymous):

thank you!